how fast and electron moving iforward it has a wavelength equal to distance it travels in one second.
 

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

If v is the speed of particle, distance travelled in 1 second= vt=v

Hence\lambda=v\implies \dfrac{h}{mv}=v
\implies v=\sqrt{\dfrac{h}{m}}